OpenStudy (ibrafacts):

Question 11. A sphere has radius 42 m. If the radius is divided by 6, what is the effect on the surface area? A. The surface area is multiplied by 36. B. The surface area is divided by 6. C. The surface area is divided by 36. D. The surface area is multiplied by 6.

OpenStudy (ibrafacts):

help me @jdoe0001

OpenStudy (jdoe0001):

hmm one sec

OpenStudy (isaidavila):

The área is divided by 36 sine the surface área of a sphere is A=4πr^2 so if r=r/6 4π(r/6)^2 4πr^2/36 that is A/36

OpenStudy (ibrafacts):

yup

OpenStudy (jdoe0001):

\(\textit{surface area of a sphere}=s={\color{blue}{ 4\pi r^2 }} \\ \quad \\ \textit{now if }r=\frac{r}{6}\quad then\implies s=4\pi \left( \cfrac{r}{6} \right)^2 \\ \quad \\ s=4\pi \left( \cfrac{r^2}{6^2} \right)\implies s=4\pi \cdot \cfrac{r^2}{6^2}\implies s=\cfrac{{\color{blue}{ 4\pi r^2 }}}{36}\)

OpenStudy (ibrafacts):

so i would divide that and i would get my answer

OpenStudy (jdoe0001):

yeap

OpenStudy (ibrafacts):

i got 12.36

OpenStudy (ibrafacts):

makes no sense

OpenStudy (jdoe0001):

well... the question isjust what happens to the surface so... notice the original, and the r/6 radius the r/6 radius one, is a fraction of the original

OpenStudy (ibrafacts):

am i suppose to get actually solve it or just find our which number i am suppose to divide with?

OpenStudy (jdoe0001):

well, based on the given choices, you're just supposed to tell, what happens to the surface area, is all and as you can see, it became 1/36 of the original :)

OpenStudy (ibrafacts):

oh its c

OpenStudy (jdoe0001):

ahemm.... yes

OpenStudy (ibrafacts):

thank you @jdoe0001 the legend

OpenStudy (jdoe0001):

yw

OpenStudy (ibrafacts):

@jdoe0001 forgot to give u a medal